OpenStudy (bmk614):

What is the area found under the curve of y=sin(x/2), x>0.

OpenStudy (bmk614):

A) 2 B) pi C) 4 D) 2pi-1 E) 2pi

OpenStudy (mathmale):

For just one cycle of the sine function? or? While not strictly required, graphing this function and establishing the beginning and ending points of x for the area to be calculated might prove to be helpful to you. Which methods have you learned for finding the area under a curve? Have you studied definite integrals? Have you estimated such areas using the Trapezoidal Rule or Simpson's Rule?

OpenStudy (bmk614):

I have learned everything In AP Calculus AB... I am reviewing for the AP exam next week.

OpenStudy (bmk614):

The the interval is between 0 and the first x-intercept

Nnesha (nnesha):

\[area =\int\limits_{a}^{b} f(x) dx\] hmm so first x-intercept at 2pi [0, 2pi] \[\int\limits_{0}^{2\pi}\sin(\frac{x}{2})dx\]e.e

Nnesha (nnesha):

take the anti-derivative and then find F(b)-F(a) here F represents the antiderivative

OpenStudy (bmk614):

Thank you! I got the right answer!

Nnesha (nnesha):

yw! :)